The average grade of two classes is 88. One class has 30 students with an average grade of 84. If the average of the second class is 96, how many students does the second class have?
Answer by oberobic(2304)

The arithmetic average = total of all students' scores / number of students.
avg = total/count
so
count*avg = total
.
We do not know the count in the second class, but we know the average, so we can use the average.
The count in the second class = x.
So the combined count in the two classes = 30+x.
.
30*84 + x*96 = (30+x)*88
.
2520 + 96x = 2640 +88x
8x = 120
x = 15
There are 15 students in the second class.
.
Done.